You have an absolutely brilliant child who is six years old and will be attending a private college in twelve years. You know that a four-year college now costs at least $30,000 per year, including tuition, books, and room and board. The cost of sending a child to college has increased by 7 percent, and you believe this will be true for the next twelve years. What will a four-year college education cost when your child is eighteen

a. $180,087.64

b. $241,463.56

c. $265,281.76

d. $270,263.00
Answer #1

It costs 30,000 per year, for 4 years the cost will be 30000*4 = 120,000

As,  A = P(1+r/100)^n

A = 120000*(1+0.07/100)^12

A = 120000*2.252191

A= 270263

Thus the answer is d)